# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

### v1

@@ Appendix A Proofs of Section 4
We write out $\operatorname{\textsc{gr}}\text{-}\epsilon(k)$ recursively like before, but now subtracting $\epsilon$ in each iteration. For any $k$, we have
$\operatorname{\textsc{gr}}\text{-}\epsilon(k)$ $=\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)+\operatorname{\textsc{gr}}\text{-}\epsilon_{k}$
$\geq\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)+\frac{\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)}{k}-\epsilon$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}\text{-}\epsilon(k-1)\left(1-\frac{1}{k}\right)-\epsilon$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}\text{-}\epsilon(k-2)\left(1-\frac{1}{k}\right)-\epsilon\right)\left(1-\frac{1}{k}\right)-\epsilon$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\operatorname{\textsc{gr}}\text{-}\epsilon(k-2)\left(1-\frac{1}{k}\right)^{2}$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{2}+\operatorname{\textsc{gr}}\text{-}\epsilon(k-3)\left(1-\frac{1}{k}\right)^{3}$$\ldots$
@@
$\ \ \ +\ldots+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{k-2}+\operatorname{\textsc{gr}}\text{-}\epsilon(1)\left(1-\frac{1}{k}\right)^{k-1}$
@@
which we can bound from below by $(\operatorname{\textsc{opt}}(k)-\epsilon k)\left(1-\frac{1}{\operatorname{\text{e}}}\right)$ again by taking a limit $k\to\infty$. This shows $\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\left(1-\frac{\epsilon k}{\operatorname{\textsc{opt}}(k)}\right)\left(1-\frac{1}{\operatorname{\text{e}}}\right)=\left(1-\frac{1}{\operatorname{\text{e}}}\right)+\frac{1}{\operatorname{\textsc{opt}}(k)}\left(\frac{\epsilon k}{\operatorname{\text{e}}}-\epsilon k\right)\iff\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)+\epsilon k\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}\geq 1-\frac{1}{\operatorname{\text{e}}}.$
Using $\operatorname{\textsc{gr}}(k)\geq k$, we can reformulate $1-\frac{1}{\operatorname{\text{e}}}\leq\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)+\epsilon k\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}\leq\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)+\operatorname{\textsc{gr}}\text{-}\epsilon(k)\cdot\epsilon\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}=\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)\cdot\left(1+\epsilon\left(1-\frac{1}{\operatorname{\text{e}}}\right)\right)}{\operatorname{\textsc{opt}}(k)}$. ∎
Theorem 4.4.
For any non-trivial instance, any $\delta>0,\,\gamma\in(0,1)$, and $k<t\leq m$, with probability at least $1-\delta$, Algorithm 2 is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover using
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in\mathcal{O}\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$
queries.
@@
We will compute the necessary values of $\epsilon$ and $\ell$ so that the statement is true, and we will see that they are indeed as stated in Algorithm 2. From Proposition 4.2 and Lemma 4.3, we know that if all our estimates are within $\frac{1}{2}\epsilon$ of the value that they estimate, then $\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\left(1-\frac{1}{\operatorname{\text{e}}}\right)\left(\frac{1}{1+\epsilon(1-\frac{1}{\operatorname{\text{e}}})}\right)$. For $\gamma\in(0,1)$, we take $\epsilon=\frac{(1-\gamma)e}{\gamma(e-1)}$ so that $\frac{\operatorname{\textsc{gr}}\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\gamma\cdot\left(1-\frac{1}{\operatorname{\text{e}}}\right)$.
We want to find $\ell$ so that with probability $1-\delta$, all our estimations $\hat{\Delta}$ are within $\frac{1}{2}\epsilon$ of the true values $\Delta$. For this, we use Hoeffding’s inequality to bound the probability that a single estimate is more than $\frac{1}{2}\epsilon$ away from its corresponding true value. We assume our samples are independent, and we know that all samples are in $\{0,1\}$, that their sample mean is $\hat{\Delta}$ and their expected value is $\Delta$, so using Hoeffding’s inequality we obtain for any estimate $\hat{\Delta}$ of $\Delta$:
@@
There are $k$ iterations and at most $m$ estimates per iteration so no more than $k\cdot m$ estimations in total. Thus, by independence, we have
@@
This means, for any $0<\delta\leq 1$, any $\ell$ that is at least this large, will, with probability $1-\delta$, produce (querying $\ell$ voters) estimates that are all at most $\frac{1}{2}\epsilon$ apart from the value that they estimate. Then we can invoke Lemma 4.3 to obtain the desired ratio. We take the smallest possible value: $\ell=\left\lceil\frac{2}{\epsilon^{2}}\log\left(\frac{2mk}{\delta}\right)\right\rceil$ and filling in for $\epsilon$ this results in
@@
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\ell\leq k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in\mathcal{O}\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$ (14)
queries. ∎
Proposition 4.9.
For any instance $(V,C,A,k)$ and any matroid $\mathcal{M}$ of rank $k$, no cover of size $\ell<k$ will attain a unique maximal Max Cover score.
@@
Suppose there does exist a cover $L\in\mathcal{M}$ of size $\ell<k$ that has a higher Max Cover score than any size-$k$ cover $M\in\mathcal{M}$. This means there exists no superset of $L$ in $\mathcal{M}$, in particular not any of size $k$, because the Max Cover score is monotone in the amount of sets that the cover contains. But by requirement 3) in Definition 4.6, using any basis element $K\in\mathcal{M}$, we can make a path of sets $L,\ L\cup k_{1},\ L\cup k_{1}\cup k_{2},\ \ldots,L\cup k_{1}\cup\ldots\cup k_{k-\ell}=:L^{\prime}$ from $L$ to attain a basis element $L^{\prime}\in\mathcal{M}$ that contains $L$. ∎
Theorem 4.10.
For any $\delta>0$, $\gamma\in(0,1)$, $m\geq t>k$, and some universal constant $C_{2}$, with probability at least $1-\delta$, Algorithm 4, with $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$, is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover when using
$\xi\alpha_{k}\cdot\left(\frac{\gamma k\log k}{C_{2}(1-\gamma)}\right)^{3}\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil\frac{(2-2/\operatorname{\text{e}})^{2}}{2}\left(\frac{2\xi}{\xi-1}\right)^{2}\log\left(\frac{2\cdot(m-k)k\cdot\xi\alpha_{k}\cdot\gamma k\log k}{C_{2}(1-\gamma)\delta}\right)\right\rceil$
queries. Asymptotically in $k,m,\gamma$ and $\delta$, this is in
$\mathcal{O}\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right).$
Proof.
We will compute the necessary values of $\epsilon$ and $\ell$ so that the statement is true, and we will see that they are indeed the same as stated in Algorithm 4. In the original approximation ratio (of Algorithm 3), we substitute $\gamma$ for $(1-1/\operatorname{\text{e}})(1-\gamma)$, so that the ratio becomes $(1-1/\operatorname{\text{e}})\cdot\gamma$ and we adapt the constant factor accordingly.
With this value of $\beta$, the minimum step size of Local Search-$\beta$ (Algorithm 3) is $C_{2}\frac{1-\gamma}{\gamma k\log k}$, and every committee attains score at least 00 and at most $\alpha_{k}$, so Algorithm 3 runs at most $\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ iterations. We anticipate that querying will increase the amount of required iterations, because of the uncertainty inducing estimations. Depending on how many iterations we allow, values of $\epsilon$ (the estimation margin) and $\ell$ (the amount of voters queried per round) will also differ. Thus, to make this dependency explicit, we add in a factor $\xi\geq 1$ and say the algorithm with queries has a total of at most $\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ iterations. We then compute the maximal value of $\epsilon$ this corresponds to.
Until the last run of the algorithm, we have $\hat{\Delta}(W,c^{\prime},c)>\beta-\epsilon$, and when this is no longer true, the swap of $c$ for $c^{\prime}$ is not made. We take $\epsilon$ so that all improvements ${\Delta}(W,c^{\prime},c)$ are at least size $\hat{\Delta}(W,c^{\prime},c)-\epsilon>\beta-2\epsilon$ and solve
@@
for $\epsilon$ to obtain $\epsilon=\frac{\xi-1}{2\xi}\cdot C_{2}\frac{1-\gamma}{\gamma k\log k}$, which is the maximum margin of estimation we can allow in order to have our original number of iterations multiplied by at most $\xi$.
We want to have $\ell$ such that with probability $1-\delta$, all our estimations $\hat{\Delta}$ are within $\epsilon$ of the true values $\Delta$. To this end, we use Hoeffding’s inequality to bound the probability that a single estimation is more than $\epsilon$ away from its corresponding $\Delta$-value. For any committee, any voter will attain some score in $\{\alpha_{0},\ldots,\alpha_{k}\}$ and whenever a candidate $c\in W$ is replaced for another candidate $c^{\prime}\notin W$, the voter can approve one less candidate than before, the same amount of candidates as before, or one more candidate than before. A score $\alpha_{i}$ before a swap, can thus become any of scores $\{\alpha_{i-1},\alpha_{i},\alpha_{i+1}\}$ after. In Lemma 3, [12] prove that for all $i<k$, $\alpha_{i+1}>\alpha_{i}$ and $\alpha_{i+2}-\alpha_{i+1}\leq\alpha_{i+1}-\alpha_{i}$. Thus, $\{\alpha_{i}\}_{i}$ is an increasing sequence, but the increase declines monotonically. We may thus take $\alpha_{1}-\alpha_{0}=1-\frac{1}{\operatorname{\text{e}}}$ for an upper bound on the difference between consecutive values in $\{\alpha_{i}\}_{i}$. Then we see that all our samples of $\Delta(W,c)$ and $\Delta(W,c^{\prime},c)$ take values in the interval $[-1+\frac{1}{\operatorname{\text{e}}},1-\frac{1}{\operatorname{\text{e}}}]$. The sample mean is $\hat{\Delta}$ and the expected value is $\Delta$, and we assume the samples are independent. Hoeffding’s inequality then yields, for any of the estimates $\hat{\Delta}$ of $\Delta$,
@@
This means, for any $0<\delta\leq 1$, any $\ell$ that is at least this size, will, with probability $1-\delta$, produce (querying $\ell$ voters) estimates that are all at most $\epsilon$ apart from the value that they estimate. We take the smallest possible value of $\ell$:
@@
The algorithm terminates when $\hat{\Delta}(W,c^{\prime},c)<\frac{C_{2}(1-\gamma)}{\gamma k\log k}-\epsilon$, which implies $\Delta(W,c^{\prime},c)<\frac{C_{2}(1-\gamma)}{\gamma k\log k}$. This is the original stopping condition of the algorithm without queries, so the original performance guarantees also hold here; the approximation ratio is $(1-1/\operatorname{\text{e}})\gamma$. Last we study the amount of queries. In total there are
@@
queries. Asymptotically in $k\ll m,\,\gamma$ and $\delta$, this is in $\mathcal{O}\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right)$. ∎
Appendix B Lower bound for the query complexity of non-adaptive algorithms in the incomplete information setting
For the incomplete information setting, we have seen two algorithms. Both algorithms make use of information acquired during the run of the algorithm. Indeed, when selecting the query sets, we need these to contain the currently elected committee. We next present a lower bound for the query complexity of non-adaptive algorithms, which can not adapt their querying strategy to readily obtained information. The proof concerns the original setting, without matroid constraints, but the result extends to the problem over a matroid constraint, as it is a more general problem. The proof strategy follows [15, Theorem 3.1].
Theorem B.1.
For any $k\geq 2$, $\epsilon>0$, and $\gamma\in(\frac{1}{2},1)$, any non-adaptive committee selection algorithm that makes fewer than $\Omega(\frac{m^{2}\epsilon}{1-\gamma})$ queries, attains a $(1-1/\operatorname{\text{e}})\gamma$-approximate max cover score with probability at most $\epsilon$.

### v2

@@ Appendix A A   Proof of Theorem 2
We write out $\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)$ recursively like before, but now subtracting $\epsilon$ in each iteration. For any $k$, we have
$\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)$ $=\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-1)+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon_{k}$
$\geq\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-1)+\frac{\operatorname{\textsc{opt}}(k)-\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-1)}{k}-\epsilon$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-1)\left(1-\frac{1}{k}\right)-\epsilon$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-2)\left(1-\frac{1}{k}\right)-\epsilon\right)\left(1-\frac{1}{k}\right)-\epsilon$
$=\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-2)\left(1-\frac{1}{k}\right)^{2}$
$\geq\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{2}+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k-3)\left(1-\frac{1}{k}\right)^{3}$ $\ldots$
@@
$\ \ \ +\ldots+\left(\frac{\operatorname{\textsc{opt}}(k)}{k}-\epsilon\right)\left(1-\frac{1}{k}\right)^{k-2}+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(1)\left(1-\frac{1}{k}\right)^{k-1}$
@@
which we can bound from below by $(\operatorname{\textsc{opt}}(k)-\epsilon k)\left(1-\frac{1}{\operatorname{\text{e}}}\right)$ again by taking a limit $k\to\infty$. This shows $\frac{\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\left(1-\frac{\epsilon k}{\operatorname{\textsc{opt}}(k)}\right)\left(1-\frac{1}{\operatorname{\text{e}}}\right)=\left(1-\frac{1}{\operatorname{\text{e}}}\right)+\frac{1}{\operatorname{\textsc{opt}}(k)}\left(\frac{\epsilon k}{\operatorname{\text{e}}}-\epsilon k\right)\iff\frac{\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)+\epsilon k\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}\geq 1-\frac{1}{\operatorname{\text{e}}}.$ Using $\operatorname{\textsc{gr}}(k)\geq k$ (Proposition 12), we can reformulate $1-\frac{1}{\operatorname{\text{e}}}\leq\frac{\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)+\epsilon k\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}\leq\frac{\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)+\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)\cdot\epsilon\left(1-\frac{1}{\operatorname{\text{e}}}\right)}{\operatorname{\textsc{opt}}(k)}=\frac{\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)\cdot\left(1+\epsilon\left(1-\frac{1}{\operatorname{\text{e}}}\right)\right)}{\operatorname{\textsc{opt}}(k)}$. ∎
Output: Committee $W\subseteq C$ of size $k$.
Let $W=\{\}$ be an empty set;
for $i=1,\ldots,k$, do
  1. 1.
    For all $c\notin W$, determine $\Delta(W,c)$.
  2. 2.
    Add to $W$ any $c^{\prime}\notin W$ for which $\Delta(W,c^{\prime})\in[\max_{c\notin W}\,\Delta(W,c)-\epsilon,\max_{c\notin W}\,\Delta(W,c)$], chosen uniformly at random.
   end for
ALGORITHM 5 A greedy-$\epsilon$ algorithm
Theorem 2.
Let an instance be non-trivial, $\delta>0,\,\gamma\in(0,1)$, and $k<t\leq m$. Then, w.p. at least $1-\delta$, Algorithm 2 is $(1-1/e)\gamma$-approximate for Max Cover with query complexity $O\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)\in O_{\delta,\gamma,k}(m\log m)$.
@@
We show that the greedy query algorithm, in order to be $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover with stated probability and number of queries, indeed needs the values of $\epsilon$ and $\ell$ as stated in Algorithm 2.
For $\gamma\in(0,1)$, we take $\epsilon=\frac{(1-\gamma)e}{\gamma(e-1)}$ so that $\frac{\operatorname{\textsc{gr}}\hskip-1.42262pt\text{-}\epsilon(k)}{\operatorname{\textsc{opt}}(k)}\geq\gamma\cdot\left(1-\frac{1}{\operatorname{\text{e}}}\right)$, using Lemma 13. We want to find $\ell$ so that with probability $1-\delta$, all our estimations $\hat{\Delta}$ are within $\frac{1}{2}\epsilon$ of the true values $\Delta$. For this, we use Hoeffding’s inequality to bound the probability that a single estimate $\hat{\Delta}$ is more than $\frac{1}{2}\epsilon$ away from its corresponding true value $\Delta$. Write $X_{W,c}:V\to\{0,1\}:v_{i}\mapsto\mathds{1}_{\{A(i)\cap(W\cup\{c\})\neq\emptyset\}}(i)-\mathds{1}_{\{A(i)\cap W\neq\emptyset\}}(i)$ for a random variable with expected value $\mathbb{E}[X_{W,c}]=\Delta(W,c)$. We can draw from $X_{W,c}$ by querying individual voters on $W\cup\{c\}$, then $\ell$ independent realizations with values in $\{0,1\}$ yield sample mean $\hat{\Delta}(W,c)$, with $\mathbb{E}[\hat{\Delta(W,c)}]=\Delta(W,c)$. Hoeffding’s inequality gives, for any estimate $\hat{\Delta}$ of $\Delta$:
@@
There are $k$ iterations and at most $m$ estimates $\hat{\Delta(W,c)}$ per iteration so no more than $k\cdot m$ estimations in total. Thus, by independence, we have
@@
This means, for any $\delta>0$, any $\ell$ that is at least this large, will, with probability $1-\delta$, produce (querying $\ell$ voters) estimates that are all at most $\frac{1}{2}\epsilon$ apart from the value that they estimate. Then we can invoke Lemma 13 to obtain the desired ratio. We take the smallest possible value: $\ell=\left\lceil\frac{2}{\epsilon^{2}}\log\left(\frac{2mk}{\delta}\right)\right\rceil$ and filling in for $\epsilon$ this results in
@@
$k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\ell\leq k\cdot t\left\lceil\frac{m-k}{t-k}\right\rceil\cdot\left\lceil 0.8\left(\frac{\gamma}{1-\gamma}\right)^{2}\log\left(\frac{2mk}{\delta}\right)\right\rceil\in O\left(\left(\frac{\gamma}{1-\gamma}\right)^{2}km\log\left(\frac{km}{\delta}\right)\right)$ (11)
queries to $A$ in total. ∎
Appendix B B   Proof of Theorem 6
We prove that ls-$\beta$-incomplete (Algorithm 4) can get arbitrarily close to the optimal approximation ratio for Max Cover using $O(m\log m)$ queries to voters. This result was stated as Theorem 6. First, we discuss some changes we made to the complete local search algorithm of [11] (resulting in Algorithm 3), as these changes shine through in the incomplete local search algorithm.
The first change we made is that we have left out the greedy initialization. Although it speeds up the convergence time-wise, it would cost many extra querying rounds. It is not possible to query only for the initialization and use the gathered information for the local search phase, since candidates not elected in the initialization, or only elected in its last round, will not have query responses from all voters whereas that is necessary if the candidate would become part of the tentative solution at some point during the local search. However, it might be possible to save the information gathered in the initialization in a smart way. For now, we leave out this phase. For similar reasons, we omit the partial enumeration step that would yield the clean $1-1/\operatorname{\text{e}}$ ratio. Secondly, we have adapted the stopping condition from $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\beta n$ (the factor $n$ is added because our scores range from $[0,1]$ instead of $[0,n]$) to $f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta$. Before we can make any kind of statement about the performance of our local search querying algorithm, we need guarantees on the accuracy of our estimations $\hat{\Delta}$. We use Hoeffding’s inequality to bound the probability that our estimates $\hat{\Delta}$ are far from the estimated true value $\Delta$, but for that we need our samples to be bounded as well. For our stopping condition $\hat{\Delta}(W,c^{\prime},c)=\hat{f}(W\cup\{c^{\prime}\}\setminus\{c\})-\hat{f}(W)<\beta-\epsilon$, these samples are really just any difference between two consecutive values of $(\alpha_{i})_{i\in\mathbb{N}}$, since replacing $c$ by $c^{\prime}$, any voter can be represented one time less, the same number of times, or one time more. The samples are thus bounded from above by $\alpha_{1}-\alpha_{0}=\alpha_{1}$. However, since $\alpha_{0}=0$, samples for the stopping condition $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\beta n$ are unbounded. Luckily, when $f((W\cup c^{\prime})\setminus c)-f(W)\leq\beta$, then also $\frac{f((W\cup c^{\prime})\setminus c)}{f(W)}\leq 1+\frac{\beta}{f(W)}$, which is at most $1+\beta n$ iff $f(W)\geq 1/n$. This will always be true as long as $\forall c_{j}\ \exists i:c_{j}\in A(i)$; all candidates have at least one approval, and $k\geq 3$, so that for any elected committee $W$, $f(W)\geq\alpha_{3}/n>1/n$.
Theorem 6.
Let $|c_{j}|\geq 1\ \forall j$, $\delta>0$, $\gamma\in(0,1)$, $m\geq t>k\geq 3$. Fix $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$ for some constant $C_{2}$. Then, w.p. at least $1-\delta$, Algorithm 4 is $(1-1/e)\gamma$-approximate for Max Cover with query complexity $O\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right)\in O_{\delta,\gamma,k}(m\log m)$.
@@ Appendix B B   Proof of Theorem 6
We show that the local search query algorithm, in order to be $(1-1/\operatorname{\text{e}})\gamma)$-approximate for Max Cover with stated probability and number of queries, indeed needs the values of $\epsilon$ and $\ell$ as in Algorithm 4. In the original approximation ratio (of Algorithm 3), we substitute $\gamma$ for $(1-1/\operatorname{\text{e}})(1-\gamma)$, so that the ratio becomes $(1-1/\operatorname{\text{e}})\cdot\gamma$ and we adapt the constant factor accordingly.
With $\beta=C_{2}\frac{1-\gamma}{\gamma k\log k}$, the minimum step size of local search-$\beta$ (Algorithm 3) is $C_{2}\frac{1-\gamma}{\gamma k\log k}$, and every committee attains score at least 0 and at most $\alpha_{k}$, so Algorithm 3 runs at most $\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ iterations. We anticipate that querying will increase the amount of required iterations, because of the uncertainty inducing estimations. Depending on how many iterations we allow, values of $\epsilon$ (the estimation margin) and $\ell$ (the number of voters queried per round) will also differ. Thus, to make this dependency explicit, we add in a factor $\xi\geq 1$ and let Algorithm 4 have a total of at most $\xi\alpha_{k}\cdot\frac{\gamma k\log k}{C_{2}(1-\gamma)}$ iterations. We then compute the maximal value of $\epsilon$ this corresponds to.
Until the last run of the algorithm, we have $\hat{\Delta}(W,c^{\prime},c)>\beta-\epsilon$, and when this is no longer true, the swap of $c$ for $c^{\prime}$ is not made. We take $\epsilon$ so that all true improvements ${\Delta}(W,c^{\prime},c)$ are at least size $\hat{\Delta}(W,c^{\prime},c)-\epsilon>\beta-2\epsilon$ and for that, we solve
@@
for $\epsilon$ to obtain $\epsilon=\frac{\xi-1}{2\xi}\cdot C_{2}\frac{1-\gamma}{\gamma k\log k}$, which is the maximum value of $\epsilon$ we may allow in order to have our original number of iterations (for Algorithm 3) multiplied by at most $\xi$.
We take $\ell$ such that with probability $1-\delta$, all our estimations $\hat{\Delta}$ are within $\epsilon$ of the true values $\Delta$. For this, we use Hoeffding’s inequality to bound the probability that a single estimate $\hat{\Delta}$ is more than $\epsilon$ away from its corresponding true value $\Delta$. Write $X_{W,c^{\prime},c}:V\to[-\alpha_{1},\alpha_{1}]:v_{i}\mapsto\alpha_{h_{i}((W\cup\{c^{\prime}\})\setminus\{c\})}-\alpha_{h_{i}(W)}$ for a random variable with expected value $\mathbb{E}[X_{W,c^{\prime},c}]=\Delta(W,c^{\prime},c)$. We can draw from $X_{W,c,c^{\prime}}$, indirectly, by querying individual voters on $(W\cup\{c^{\prime}\})\setminus\{c\}$, then $\ell$ independent realizations yield sample mean $\hat{\Delta}(W,c^{\prime},c)$, with $\mathbb{E}[\hat{\Delta(W,c^{\prime},c)}]=\Delta(W,c^{\prime},c)$. Regarding the range of $X_{W,c^{\prime},c}$: For any committee $W$, any voter $v_{i}$ will attain some score $\alpha_{h_{i}(W)}\in\{\alpha_{0},\ldots,\alpha_{k}\}$ and whenever a candidate $c\in W$ is replaced for another candidate $c^{\prime}\notin W$, the voter can approve one less candidate than before, the same number of candidates as before, or one more candidate than before. A score $\alpha_{i}$ before a swap, can thus become any of scores $\{\alpha_{i-1},\alpha_{i},\alpha_{i+1}\}$ after. In Lemma 3, [11] prove that for all $i<k$, $\alpha_{i+1}>\alpha_{i}$ and $\alpha_{i+2}-\alpha_{i+1}\leq\alpha_{i+1}-\alpha_{i}$. Thus, $\{\alpha_{i}\}_{i}$ is an increasing sequence, but the increase declines monotonically. We may thus take $\alpha_{1}-\alpha_{0}=1-\frac{1}{\operatorname{\text{e}}}$ for an upper bound on the difference between consecutive values in $\{\alpha_{i}\}_{i}$. Then we see that all our samples of $X_{W,c^{\prime},c}$ take values in the interval $1+\frac{1}{\operatorname{\text{e}}},1-\frac{1}{\operatorname{\text{e}}}]$. Combining the above, Hoeffding’s inequality gives, for any estimate $\hat{\Delta}$ of $\Delta$:
@@
This means, for any $\delta>0$, any $\ell$ that is at least this size, will, with probability $1-\delta$, produce (querying $\ell$ voters) estimates that are all at most $\epsilon$ apart from the value that they estimate. We take the smallest possible value of $\ell$:
@@
The algorithm terminates when $\hat{\Delta}(W,c^{\prime},c)<\frac{C_{2}(1-\gamma)}{\gamma k\log k}-\epsilon$, which implies $\Delta(W,c^{\prime},c)<\frac{C_{2}(1-\gamma)}{\gamma k\log k}$. This is the original stopping condition of the algorithm without queries, so the original performance guarantees also hold here; the approximation ratio is $(1-1/\operatorname{\text{e}})\gamma$. Last we study the number of queries. In total there are
@@
queries to $A$. Asymptotically in $\gamma,\ \delta,\ k$ and $m$ (recall $k\ll m$), this is in $O\left(\left(\frac{\gamma k\log k}{1-\gamma}\right)^{3}m\log\left(\frac{m\gamma k^{2}\log k}{\delta(1-\gamma)}\right)\right)$. ∎
Appendix C C   Lower bound for non-adaptive algorithms in the incomplete information setting
For the incomplete information setting, we have seen two algorithms. Both algorithms make use of information acquired during the run of the algorithm. Indeed, when selecting the query sets, we need these to contain the currently elected committee. We next present a lower bound for the query complexity of non-adaptive algorithms, which can not adapt their querying strategy to readily obtained information. The proof concerns the original setting, without matroid constraints, but the result extends to the problem over a matroid constraint, as it is a more general problem.
Theorem 14.
For any $k\geq 2$, $\epsilon>0$, and $\gamma\in(\frac{1}{2},1)$, any non-adaptive committee selection algorithm that makes fewer than $\Omega(\frac{m^{2}\epsilon}{1-\gamma})$ queries, is $(1-1/\operatorname{\text{e}})\gamma$-approximate for Max Cover with probability at most $\epsilon$.